A charter company will provide a plane for a fare of $200 each for 80 or fewer passengers. For each passenger in excess of 80, the fare is decreased by $2.00 per person for everyone. What number of passengers would produce the greatest revenue for the company.
Answer by stanbon(75887)

A charter company will provide a plane for a fare of $200 each for 80 or fewer passengers.
For each passenger in excess of 80, the fare is decreased by $2.00 per person for everyone.
What number of passengers would produce the greatest revenue for the company.
------------------
Equation:
Revenue(x) = (80+x)(200-2x)
R(x) = -2x^2 + 40x + 16000
---
Maximum Revenue occurs when x = -b/2a = -40/(-4) = 10
---------------------
80+10 = 90 would be the number of passengers needed to maximize Revenue.
============================================================================
Cheers,
Stan H.